# MME: A Comprehensive Evaluation Benchmark for Multimodal Large Language Models

### v1

@@ Abstract
Multimodal Large Language Model (MLLM) relies on the powerful LLM to perform multimodal tasks, showing amazing emergent abilities in recent studies, such as writing poems based on an image. However, it is difficult for these case studies to fully reflect the performance of MLLM, lacking a comprehensive evaluation. In this paper, we fill in this blank, presenting the first MLLM Evaluation benchmark MME [^1] Please email guilinli@stu.xmu.edu.cn to obtain the benchmark dataset. This is the V1 version of MME, which will be updated as MLLMs evolve. An online leaderboard will be released at https://github.com/BradyFU/Awesome-Multimodal-Large-Language-Models.. It measures both perception and cognition abilities on a total of 14 subtasks. In order to avoid data leakage that may arise from direct use of public datasets for evaluation, the annotations of instruction-answer pairs are all manually designed. The concise instruction design allows us to fairly compare MLLMs, instead of struggling in prompt engineering. Besides, with such an instruction, we can also easily carry out quantitative statistics. A total of 10 advanced MLLMs are comprehensively evaluated on our MME, which not only suggests that existing MLLMs still have a large room for improvement, but also reveals the potential directions for the subsequent model optimization.
@@ 1 Introduction
The thriving of Large Language Model (LLM) has paved a new road to the multimodal field, i.e., Multimodal Large Language Model (MLLM) OpenAI (2023); Alayrac et al. (2022); Huang et al. (2023); Li et al. (2023c); Driess et al. (2023). It refers to using LLM as a brain to process multimodal information and give reasoning results. Equipped with the powerful LLM, MLLM is expected to address more complex multi-modal tasks Driess et al. (2023); Wu et al. (2023); Shen et al. (2023). The three representative abilities of LLM Zhao et al. (2023a), including In-Context Learning (ICL) Brown et al. (2020), instruction following Touvron et al. (2023), and Chain-of-Thought (CoT) Wei et al. (2022), are also manifested in multimodality. For example, Flamingo Alayrac et al. (2022) turns on multimodal ICL, which can adapt to new tasks by giving a few examples. MiniGPT-4 Zhu et al. (2023) implements GPT-4OpenAI (2023)-like instruction following capabilities, such as converting images into corresponding website codes, by introducing multimodal instruction tuning. PaLM-E Driess et al. (2023) achieves amazing OCR-free math reasoning via CoT. These emergent abilities of MLLM are exciting and imply that a new dawn has broken in artificial intelligence.
Although these models exhibit surprising conversational capabilities when conducting everyday chats, we still know little about how well they quantitatively perform in various aspects. The existing three common quantitative evaluation manners for MLLMs have their limitations that are difficult to comprehensively evaluate performance. Specifically, the first manner Xu et al. (2022); Dai et al. (2023); Wang et al. (2023) evaluates on existing traditional multimodal datasets, such as image caption Chen et al. (2015) and VQA Goyal et al. (2017); Marino et al. (2019); Lu et al. (2022). However, on the one hand, it may be hard to reflect the emergent abilities of MLLM on these datasets. On the other hand, since the training sets of large models are no longer unified, it is difficult to guarantee that all MLLMs have not used the testing set for training. The second manner Ye et al. (2023) is to collect data for an open-ended evaluation, but either the data is unavailable to public by now Zhao et al. (2023c) or the amount is small (only 50 images) Ye et al. (2023). The third manner focuses on one aspect of MLLMs, such as object hallucination Li et al. (2023d) or adversarial robustness Zhao et al. (2023b), which is powerless to comprehensive evaluation.
In light of these concerns, a new comprehensive evaluation benchmark is urgently needed to match the flourish of MLLMs. We argue that a universal comprehensive evaluation benchmark should have the following four characteristics: (1) It should cover as much as possible, including both perception and cognition abilities. The former refers to recognizing the specific object, such as its existence, count, position, and color. The latter refers to compositing the perception information and the knowledge in LLM to deduce more complex answers. It is obvious that the former is the premise of the latter. (2) Its data or annotations should not come from existing publicly available datasets as much as possible, avoiding the risk of data leakage. (3) Its instructions should be as concise as possible and in line with human cognition. Although instruction design may have a large impact on the output, all models should be tested under the same unified instructions for fair comparison. A good MLLM should be able to generalize to such concise instructions. (4) The responses of MLLM to the instructions should be intuitive and convenient for quantitative analysis. The open-ended answer of MLLM poses significant challenges to the quantization. Existing methods tend to use GPT or manual scoring Li et al. (2023a); Liu et al. (2023); Ye et al. (2023), but there may be problems of inaccuracy and subjectivity.
@@
Figure 1: Diagram of our MME benchmark. It evaluates MLLM from both perception and cognition, including a total of 14 subtasks. Each image corresponds to two questions whose answers are marked yes [Y] and no [N], respectively. The instruction consists of a question followed by "Please answer yes or no". It is worth noting that all instructions are manually designed.
@@
We conduct massive experiments to evaluate the zero-shot performance of 10 advanced MLLMs on 14 subtasks. The evaluated MLLMs include BLIP-2 Li et al. (2023c), LLaVA Liu et al. (2023), MiniGPT-4 Zhu et al. (2023), mPLUG-Owl Ye et al. (2023), LLaMA-Adapter-v2 Gao et al. (2023), Otter Li et al. (2023b), Multimodal-GPT Gong et al. (2023), InstructBLIP Dai et al. (2023), VisualGLM-6B git (2023i), and PandaGPT Su et al. (2023). As displayed in Fig. 2 that consists of 2 overall leaderboards (perception and cognition) and 14 individual leaderboards, these MLLMs show clear discrepancies in our MME evaluation benchmark. Fig. 3 also provides a comparison from the other perspective. More importantly, we have summarized four prominent problems exposed in experiments, including inability to follow basic instructions, a lack of basic perception and reasoning, as well as object hallucination, as shown in Fig. 4. These findings are instructive for the subsequent model optimization.
@@
In summary, the contributions of this work are as follows: (1) We propose a new benchmark MME to meet the urgent need of MLLM evaluation. As far as we know, MME is the first comprehensive MLLM evaluation benchmark [^2] The concurrent works are LAMM-Benchmark Yin et al. (2023) and LVLM-eHub Xu et al. (2023), which are rather different from our MME. We recommend reading the two works as well to enrich the understanding of MLLM evaluation.. (2) We conduct abundant experiments on a total of 14 subtasks to comprehensively evaluate 10 up-to-date MLLMs. (3) We summarize the exposed problems in experiments, proving guidance for the evolution of MLLM.
@@ 2.1 Instruction Design
In order to facilitate quantitative performance statistics, the orientation of our instruction design is to let the model to answer “yes” or “no”. As a result, the instruction consists of two parts, including a concise question and a description “Please answer yes or no.” For each test image, we manually design two instructions, where the discrepancy lies in the questions. The ground truth answer of the first question is “yes” and that of the second question is “no”, as shown in Fig. 1. When MLLM answers both of these questions correctly, it appears more confident that the MLLM actually comprehends the image and the corresponding knowledge behind it.
@@ 2.2 Evaluation Metric
Since the output of the model is limited to two types (“yes” or “no”), it is convenient to measure the metrics of accuracy and accuracy+. The former is calculated based on each question, while the latter is based on each image where both of the two questions need to be answered correctly. The random accuracies of the two metrics are equal to 50% and 25%, respectively. It can be seen that accuracy+ is a stricter measurement but also better reflects the comprehensive understanding degree of the model to the image. In addition, we calculate the score of a subtask based on the sum of accuracy and accuracy+. The perception score is the sum of scores of all perception subtasks. The cognition score is calculated in the same way.
Figure 2: Leaderboards on our MME benchmark. (1) and (2) are the overall leaderboards of perception and cognition respectively, in which the full score of the former is 2000 and that of the latter is 800. (3)-(16) are the leaderboards of the 14 subtasks with the full score of 200. The score is the sum of the accuracy and the accuracy+ in Tables. 1, 2, and 3.
Figure 3: Comparison of 10 advanced MLLMs on 14 subtasks. The full score of each subtask is 200.

### v4

@@ Abstract
Multimodal Large Language Model (MLLM) relies on the powerful LLM to perform multimodal tasks, showing amazing emergent abilities in recent studies, such as writing poems based on an image. However, it is difficult for these case studies to fully reflect the performance of MLLM, lacking a comprehensive evaluation. In this paper, we fill in this blank, presenting the first comprehensive MLLM Evaluation benchmark MME [^1] MME is collected by Xiamen University for academic research only. This is the v1 version and will be updated as MLLMs evolve.. It measures both perception and cognition abilities on a total of 14 subtasks. In order to avoid data leakage that may arise from direct use of public datasets for evaluation, the annotations of instruction-answer pairs are all manually designed. The concise instruction design allows us to fairly compare MLLMs, instead of struggling in prompt engineering. Besides, with such an instruction, we can also easily carry out quantitative statistics. A total of 30 advanced MLLMs are comprehensively evaluated on our MME, which not only suggests that existing MLLMs still have a large room for improvement, but also reveals the potential directions for the subsequent model optimization. The data application manner and online leaderboards are released at https://github.com/BradyFU/Awesome-Multimodal-Large-Language-Models/tree/Evaluation.
@@ 1 Introduction
The thriving of Large Language Model (LLM) has paved a new road to the multimodal field, i.e., Multimodal Large Language Model (MLLM) [37, 8, 20, 23, 13, 9]. It refers to using LLM as a brain to process multimodal information and give reasoning results [49]. Equipped with the powerful LLM, MLLM is expected to address more complex multi-modal tasks [13, 45, 38, 27, 54, 31]. The three representative abilities of LLM [55], including instruction following [40], In-Context Learning (ICL) [10], and Chain-of-Thought (CoT) [43] are also manifested in multimodality. For example, Flamingo [8] turns on multimodal ICL, which can adapt to new tasks by giving a few examples. PaLM-E [13] achieves amazing OCR-free math reasoning via CoT. GPT-4V [37] shows even more ability in a variety of complex reasoning tasks [46]. MiniGPT-4 [59] implements GPT-4[37]-like instruction following capabilities, such as converting images into corresponding website codes, by introducing multimodal instruction tuning. These emergent abilities of MLLMs are exciting and imply that a new dawn has broken in artificial intelligence.
@@
Figure 1: Diagram of our MME benchmark. It evaluates MLLMs from both perception and cognition, including a total of 14 subtasks. Each image corresponds to two questions whose answers are marked yes [Y] and no [N], respectively. The instruction consists of a question followed by “Please answer yes or no”. It is worth noting that all instructions are manually designed.
Figure 2: Leaderboards on our MME benchmark. (1) and (2) are the overall leaderboards of perception and cognition respectively, in which the full score of the former is 2000 and that of the latter is 800. (3)-(16) are the leaderboards of the 14 subtasks with the full score of 200. The score is the sum of the accuracy and the accuracy+ in Tables. 1 and 2. A total of 30 advanced MLLMs joint the leaderboards. For the sake of presentation, we only show 10 models for each list, in which the top three ones are given clear trophy logos.
Although these models exhibit surprising conversational capabilities when conducting everyday chats, we still know little about how well they quantitatively perform in various aspects. The existing three common quantitative evaluation manners for MLLMs have their limitations that are difficult to comprehensively evaluate performance. Specifically, the first manner [47, 12, 42] evaluates on existing traditional multimodal datasets, such as image caption [11] and VQA [16, 36, 32]. However, on the one hand, it may be hard to reflect the emergent abilities of MLLMs on these datasets. On the other hand, since the training sets of large models are no longer unified, it is difficult to guarantee that all MLLMs have not used the testing set for training. The second manner [48] is to collect data for an open-ended evaluation, but either the data is unavailable to public by now [57] or the amount is small (only 50 images) [48]. The third manner focuses on one aspect of MLLMs, such as object hallucination [25] or adversarial robustness [56], which is powerless to comprehensive evaluation.
In light of these concerns, a new comprehensive evaluation benchmark is urgently needed to match the flourish of MLLMs. We argue that a universal comprehensive evaluation benchmark should have the following four characteristics: (1) It should cover as much as possible, including both perception and cognition abilities. The former refers to recognizing the specific object, such as its existence, count, position, and color. The latter refers to compositing the perception information and the knowledge in LLM to deduce more complex answers. It is obvious that the former is the premise of the latter. (2) Its data or annotations should not come from existing publicly available datasets as much as possible, avoiding the risk of data leakage. (3) Its instructions should be as concise as possible and in line with human cognition. Although instruction design may have a large impact on the output, all models should be tested under the same unified instructions for fair comparison. A good MLLM should be able to generalize to such concise instructions. (4) The responses of MLLMs to the instructions should be intuitive and convenient for quantitative analysis. The open-ended answer of MLLMs poses significant challenges to the quantization. Existing methods tend to use GPT or manual scoring [21, 29, 48], but there may be problems of inaccuracy and subjectivity.
@@
We conduct massive experiments to evaluate the zero-shot performance of 30 advanced MLLMs on the 14 subtasks. The evaluated MLLMs include BLIP-2 [23], InstructBLIP [12], MiniGPT-4 [59], PandaGPT [39], Multimodal-GPT [15], VisualGLM-6B [5], ImageBind-LLM [17], VPGTrans [53], LaVIN [33], mPLUG-Owl [48], Octopus [3], Muffin [51], Otter [22], LRV-Instruction [28], Cheetor [24], LLaMA-Adapter-v2 [14], GIT2 [41], BLIVA [18], Lynx [52], MMICL [54], GPT-4V [37], Skywork-MM [4], mPLUG-Owl2 [48], Qwen-VL-Chat [9], XComposer-VL [7], LLaVA [29], Lion [2], SPHINX [27], InfMLLM [1], and WeMM [6]. As displayed in Fig. 2 that consists of 2 overall leaderboards (perception and cognition) and 14 individual leaderboards, these MLLMs show clear discrepancies in our MME evaluation benchmark. Fig. 3 also provides a comparison from the other perspective. We can see the range that current MLLMs can reach in each capability dimension. More importantly, we have summarized four prominent problems exposed in experiments, including inability to follow basic instructions, a lack of basic perception and reasoning, as well as object hallucination [50, 25], as shown in Fig. 4. It is expected that these findings are instructive for the subsequent model optimization.
@@
In summary, the contributions of this work are as follows: (1) We propose a new benchmark MME to meet the urgent need of MLLM evaluation. (2) A total of 30 up-to-date MLLMs are evaluated on our MME. (3) We summarize the exposed problems in experiments, proving guidance for the evolution of MLLMs.
Model Existence Count Position Color OCR Poseter Celebrity
ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+
BLIP-2 86.67 73.33 75.00 60.00 56.67 16.67 81.67 66.67 70.00 40.00 79.25 62.59 68.53 37.06
mPLUG-Owl 73.33 46.67 50.00 0.00 50.00 0.00 51.67 3.33 55.00 10.00 78.23 57.82 66.18 34.12
ImageBind-LLM 75.00 53.33 50.00 10.00 43.33 3.33 56.67 16.67 60.00 20.00 52.72 12.24 55.29 21.18
InstructBLIP 95.00 90.00 80.00 63.33 53.33 13.33 83.33 70.00 57.50 15.00 74.15 49.66 67.06 34.12
VisualGLM-6B 61.67 23.33 50.00 0.00 48.33 0.00 51.67 3.33 42.50 0.00 53.74 12.24 50.88 2.35
Multimodal-GPT 48.33 13.33 48.33 6.67 45.00 13.33 55.00 13.33 57.50 25.00 42.86 14.97 49.12 24.71
PandaGPT 56.67 13.33 50.00 0.00 50.00 0.00 50.00 0.00 50.00 0.00 56.80 19.73 46.47 10.59
LaVIN 95.00 90.00 61.67 26.67 53.33 10.00 58.33 16.67 67.50 40.00 59.18 20.41 37.94 9.41
Cheetor 93.33 86.67 63.33 33.33 56.67 23.33 70.00 46.67 65.00 35.00 81.29 65.99 87.65 76.47
GIT2 96.67 93.33 71.67 46.67 60.00 36.67 85.00 73.33 55.00 10.00 61.56 51.02 81.18 64.71
GPT-4V 96.67 93.33 86.67 73.33 65.00 30.00 80.00 70.00 95.00 90.00 96.94 95.24 0.00 0.00
XComposer-VL 96.67 93.33 85.00 73.33 73.33 53.33 88.33 76.67 75.00 50.00 85.71 76.19 83.24 67.06
LLaVA 95.00 90.00 85.00 70.00 76.67 56.67 90.00 80.00 75.00 50.00 86.39 74.15 83.53 69.41
LRV-Instruction 88.33 76.67 68.33 43.33 56.67 30.00 88.33 76.67 70.00 40.00 78.77 60.27 67.35 45.29
Lion 96.67 93.33 85.00 70.00 83.33 70.00 93.33 86.67 57.50 15.00 93.88 87.76 82.94 67.65
Lynx 98.33 96.67 81.67 70.00 60.00 30.00 90.00 80.00 57.50 20.00 74.49 50.34 71.76 46.47
MMICL 90.00 80.00 86.67 73.33 55.00 26.67 83.33 73.33 60.00 40.00 81.63 64.63 79.41 62.35
Muffin 98.33 96.67 86.67 76.67 53.33 13.33 88.33 76.67 52.50 5.00 78.57 59.18 56.47 25.29
Octopus 93.33 86.67 50.00 3.33 45.00 3.33 66.67 36.67 55.00 10.00 78.23 59.86 75.29 54.12
Otter 98.33 96.67 58.33 30.00 60.00 26.67 70.00 43.33 57.50 15.00 78.91 59.86 90.88 81.76
Qwen-VL-Chat 85.00 73.33 83.33 66.67 75.00 53.33 90.00 80.00 80.00 60.00 92.18 86.39 72.35 48.24
SPHINX 98.33 96.67 86.67 73.33 83.33 70.00 86.67 73.33 62.50 25.00 87.41 76.87 92.65 85.29
Skywork-$M$ 93.33 86.67 81.67 70.00 46.67 16.67 81.67 63.33 87.50 75.00 91.50 84.35 86.76 73.53
VPGTrans 56.67 13.33 61.67 23.33 53.33 10.00 56.67 16.67 57.50 20.00 60.88 23.13 51.18 2.35
WeMM 98.33 96.67 80.00 60.00 73.33 53.33 88.33 80.00 82.50 65.00 86.39 74.15 92.65 86.47
BLIVA 93.33 86.67 78.33 60.00 58.33 23.33 93.33 86.67 62.50 25.00 84.35 70.75 80.29 60.59
InfMLLM 96.67 93.33 81.67 70.00 80.00 63.33 95.00 90.00 77.50 55.00 87.07 76.19 86.18 75.29
LLaMA-AdapterV2 95.00 90.00 73.33 60.00 50.00 6.67 71.67 46.67 67.50 35.00 81.97 65.99 77.94 58.82
MiniGPT-4 51.67 16.67 48.33 6.67 43.33 0.00 55.00 20.00 52.50 5.00 36.39 5.44 45.00 9.41
mPLUG-Owl2 95.00 90.00 85.00 70.00 61.67 26.67 83.33 66.67 67.50 35.00 86.73 73.47 87.94 76.47
Table 1: Evaluation results on the subtasks of existence, count, position, color, OCR, poster, and celebrity. The top two results on each subtask are bolded and underlined, respectively.
Model Scene Landmark Artwork Comm. Num. Text. Code.
ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+ ACC ACC+
BLIP-2 81.25 64.00 79.00 59.00 76.50 60.00 68.57 41.43 40.00 0.00 55.00 10.00 55.00 20.00
mPLUG-Owl 78.00 57.50 86.25 73.00 63.25 33.00 57.14 21.43 50.00 10.00 60.00 20.00 47.50 10.00
ImageBind-LLM 68.75 44.50 53.00 9.00 54.25 16.50 40.00 8.57 50.00 5.00 50.00 0.00 50.00 10.00
InstructBLIP 84.00 69.00 59.75 20.00 76.75 57.50 75.00 54.29 35.00 5.00 55.00 10.00 47.50 10.00
VisualGLM-6B 81.75 64.50 59.75 24.00 55.25 20.00 35.00 4.29 45.00 0.00 50.00 0.00 47.50 0.00
Multimodal-GPT 50.50 17.50 48.25 21.50 46.50 13.00 43.57 5.71 42.50 20.00 50.00 10.00 45.00 10.00
PandaGPT 72.50 45.50 56.25 13.50 50.25 1.00 56.43 17.14 50.00 0.00 52.50 5.00 47.50 0.00
LaVIN 78.75 58.00 64.00 29.50 59.25 28.00 58.57 28.57 55.00 10.00 47.50 0.00 50.00 0.00
Cheetor 84.50 71.50 81.41 64.32 67.50 46.00 64.29 34.29 57.50 20.00 42.50 15.00 57.50 30.00
GIT2 86.00 72.50 79.50 61.00 81.25 65.00 66.43 32.86 40.00 10.00 52.50 15.00 45.00 0.00
GPT-4V 83.50 67.50 79.25 59.00 82.00 66.00 79.29 62.86 75.00 55.00 55.00 20.00 90.00 80.00
XComposer-VL 86.25 73.50 87.75 77.50 73.25 53.00 77.14 61.43 40.00 15.00 67.50 45.00 55.00 30.00
LLaVA 86.75 74.50 90.00 80.50 70.75 47.00 73.57 54.29 37.50 5.00 57.50 20.00 42.50 5.00
LRV-Instruction 81.04 65.93 86.84 73.68 63.75 37.50 65.00 35.71 45.00 25.00 55.00 30.00 57.50 15.00
Lion 85.50 73.50 91.00 82.00 75.75 55.00 74.29 51.43 65.00 40.00 82.50 65.00 42.50 25.00
Lynx 88.00 76.50 87.00 75.00 72.50 47.00 66.43 44.29 17.50 0.00 42.50 0.00 45.00 0.00
MMICL 83.75 70.00 76.96 59.16 76.50 59.00 76.43 60.00 47.50 35.00 72.50 60.00 47.50 30.00
Muffin 83.75 67.50 81.25 65.00 71.00 45.50 68.57 41.43 45.00 0.00 57.50 15.00 47.50 15.00
Octopus 84.75 72.50 75.00 51.00 64.50 30.50 64.29 35.71 47.50 0.00 62.50 40.00 47.50 15.00
Otter 86.25 72.50 78.75 58.50 75.00 54.00 66.43 40.00 52.50 20.00 52.50 5.00 55.00 15.00
Qwen-VL-Chat 83.75 68.50 88.00 76.00 73.50 52.00 76.43 54.29 35.00 5.00 82.50 65.00 32.50 10.00
SPHINX 86.50 73.50 89.20 78.89 77.00 57.00 75.71 54.29 40.00 15.00 50.00 25.00 50.00 0.00
Skywork-$M$ 79.29 59.60 75.51 51.53 69.43 45.08 72.14 54.29 65.00 30.00 60.00 20.00 45.00 10.00
VPGTrans 80.25 61.50 54.75 10.00 57.75 19.50 50.00 14.29 50.00 0.00 57.50 20.00 52.50 5.00
WeMM 91.75 84.50 90.75 81.50 85.00 71.00 80.00 60.00 47.50 10.00 75.00 55.00 72.50 45.00
BLIVA 83.50 68.00 63.00 26.50 77.25 56.00 77.86 58.57 47.50 10.00 57.50 20.00 50.00 10.00
InfMLLM 87.75 77.50 88.50 78.50 68.00 40.50 76.43 55.71 45.00 15.00 67.50 35.00 42.50 10.00
LLaMA-AdapterV2 85.25 71.00 88.44 79.40 73.25 50.50 67.86 38.57 47.50 0.00 67.50 45.00 60.00 30.00
MiniGPT-4 54.25 17.50 45.50 8.50 50.00 10.50 46.43 12.86 45.00 0.00 0.00 0.00 40.00 0.00
mPLUG-Owl2 83.25 70.00 85.75 71.50 77.25 57.00 71.43 44.29 35.00 0.00 67.50 35.00 45.00 15.00
Table 2: Evaluation results on the subtasks of scene, landmark, artwork, commonsense reasoning, numerical calculation, text translation, and code reasoning. The top two results on each subtask are bolded and underlined, respectively.
@@ 2.1 Instruction Design
In order to facilitate quantitative performance statistics, the orientation of our instruction design is to let the model to answer “yes” or “no”. As a result, the instruction consists of two parts, including a concise question and a description “Please answer yes or no.” For each test image, we manually design two instructions, where the discrepancy lies in the questions. The ground truth answer of the first question is “yes” and that of the second question is “no”, as shown in Fig. 1. When MLLM answers both of the two questions correctly, it appears more confident that the MLLM actually comprehends the image and the corresponding knowledge behind it, rather than just guessing.
@@ 2.2 Evaluation Metric
Since the output of the model is limited to two types (“yes” or “no”), it is convenient to measure the metrics of accuracy and accuracy+. The former is calculated based on each question, while the latter is based on each image where both of the two questions need to be answered correctly. The random accuracies of the two metrics are equal to 50% and 25%, respectively. It can be seen that accuracy+ is a stricter measurement but also better reflects the comprehensive understanding degree of the model to the image. In addition, we calculate the score of a subtask based on the sum of accuracy and accuracy+. The perception score is the sum of scores of all perception subtasks. The cognition score is calculated in the same way. Therefore, the full scores of perception and cognition are 2000 and 800, respectively.
Figure 3: Comparison of 6 top MLLMs on 14 subtasks. The full score of each subtask is 200.